Assuming that the central government decides to cut taxes by R100 billion to
stimulate the economy. The relevant marginal propensity to consume is 0.6 (60
percent). What will be the impact of such fiscal policy on equilibrium GDP?
1
Expert's answer
2020-04-21T11:06:55-0400

"\frac {\Delta GDP}{100}=\frac {1}{1-0.6}; 
\Delta GDP=100/0.4;"

"\Delta GDP=" $ "250" billion;

answer: GDP will increase by $ 250 billion.